OpenStudy (anonymous):

given T(e1), T(e2), and v how do i find T(v)?

OpenStudy (lgbasallote):

are T(e1) and T(e2) functiions?

OpenStudy (anonymous):

T(e1) = (1, √3) and T(e2)= (-√3, 1)

OpenStudy (amistre64):

e1 and e2 are standard vectors (1,0) and (0,1) T is a transformation matrix with column vectors, say: u1 and u2 T(e1) = 1u1 + 0u2 = (1,sqrt(3)) T(e2) = 0u1 + 1u2 = (-sqrt(3),1) I know theres a simple way to do it, but that method wasnt a strong point in linear algebra for me. T= a c b d T(e1) = 1(a) + 0(c) = (1) 1(b) 0(d) = (sqrt(3)) hence a=1, b=sqrt(3) T(e2) = 0(a) + 1(c) = (-sqrt(3)) 0(b) 1(d) = (1) hence c=-sqrt(3), d=1 \[T=\begin{pmatrix} 1&-\sqrt{3}\\\sqrt{3}&1 \end{pmatrix}\] thats how i usually did it in class